Joe348:

An apple has 6 J of energy. The apple sits on a counter 3 m high. What is the mass of the apple?

Astro:

Do you know we should go about this problem?

Astro:

@astro wrote:
Do you know we should go about this problem?
How we should go about this problem**

Joe348:

No clue

Joe348:

Ham?

Astro:

So we know that the potential energy of the apple on the counter is 6 Joules. The counter is 3.00 Meters high. The formula for potential energy is pe=mgh PE=Potential Energy M=Mass H=Height G=Acceleration of Gravity pe=6 joules the h=3.00 meters 6=mx9.8x3 simplify: 6=M*29.4 m=6/29.4 m=.2040816 Round: M=.2KG Meaning that the mass of te apple is .2KG

Astro:

It's a long process but I hope that makes sense

Joe348:

oh wait i thought it was GPE and yes I got it

Astro:

@joe348 wrote:
oh wait i thought it was GPE and yes I got it
Why would it be GPE?

Joe348:

so its just PE?

Joe348:

@astro wrote:
@joe348 wrote:
oh wait i thought it was GPE and yes I got it
Why would it be GPE?
idk i only have GPE or KE

Astro:

@joe348 wrote:
so its just PE?
I mean I would think it would be GPE, I think it is the same thing?

Astro:

You would know better than me.

Joe348:

anyway i think your right thank you, you dont mind if i copy do ya?

Astro:

No I don't mind, I would like @vocaloid to clarify though. About the GPE and PE

Joe348:

where did you get 9.8?

Astro:

I think it is GPE though because that is when something is positioned above the height of zero

Joe348:

wait your right thank you

Astro:

@joe348 wrote:
wait your right thank you
Ok, Does it make sense though?

Joe348:

Yes

Astro:

@joe348 wrote:
Yes
Awesome! Have a good night Joe~